NEXT
Make New Entry,
Make Followup Entry
User name V. Sulkosky
Log entry time 14:17:27 on November 03, 2008
Entry number 246457
keyword=Pedestal Shifts for S2m Right ADC Channels
I have noticed that the pedestals for the 16 channels on the S2m detector
right-side have been collectively moving by about 10 channels. Before
starting
up I found the pedestals had moved higher by about 10 channels. This
morning I discovered they all shifted lower by about 20 channels.
The S2m right signal cables are plugged into the 1881 ADC in ROC 3,
slot 25, channels 48-63. I have not observed this behavior for any of
the other channels in ADC. Below I compare the left and right-side
pedestals prior to the experiment and from run 2743 taken this morning.
The left-side pedestals are very stable.
Prior to the S2m installation these 16 ADC channels were unused. I
suspect
that either the ribbon cable has an issue or this set of
16 channels has some problem. However, this ribbon cable had been
used
previously when S2 was installed. Tomorrow, I will investigate
this
problem during the access.
Left-side of S2m
PMT # Pedestal (10/25/08) Pedestal (11/3/08)
=====================================================
1 297 296
2 278 278
3 243 244
4 259 260
5 429 430
6 462 464
7 411 414
8 446 449
9 311 311
10 286 287
11 245 245
12 254 255
13 484 486
14 414 416
15 465 468
16 431 434
Right-side of S2m
PMT # Pedestal (10/25/08) Pedestal (11/3/08)
=====================================================
1 380 360
2 312 292
3 303 284
4 285 264
5 436 417
6 484 464
7 455 435
8 437 417
9 339 318
10 348 329
11 270 249
12 314 293
13 490 472
14 485 465
15 473 454
16 400 380
A copy of this log entry has been emailed to: rom,camsonne,brads,chiran,ycwang